# Towards Secure Virtual Elections: Multiparty Computation of Order Based Voting Rules

### v4

@@ 3 A secure order based voting protocol
In this section, we describe our method for securely computing the winners in two order-based voting rules, Copeland and Maximin. We begin with formal definitions in Section 3.1. Then, in Section 3.2, we characterize legal ballot matrices for each of the two rules. Such a characterization is an essential part of our method since the talliers need to verify, in an oblivious manner, that each cast ballot is indeed legal, and does not hide a malicious attempt to cheat or sabotage the elections. In other words, the talliers need to verify the legality of each cast ballot only through its secret shares, without recovering the actual ballot. The characterization that we describe in Section 3.2 will be used later on to perform such an oblivious validation.
@@
Then, in Section 3.3 we introduce our secure voting protocol. That protocol description includes a sub-protocol for validating the cast ballots and sub-protocols for computing the final election results from all legal ballots. The validation sub-protocols are described in Sections 3.4 and 3.5. The sub-protocols for computing the election results are described in Sections 3.6 and 3.7 for Copeland and Maximin rules, respectively.
@@
We conclude this section with a discussion of the overall security of our protocol (Section 3.8) and its implementation in an end-to-end voting system (Section 3.9).
@@ 3.1 Formal definitions
We consider a setting in which there are $N$ voters, $\mbox{\bf V}=\{V_{1},\ldots,V_{N}\}$, that wish to hold an election over $M$ candidates, $\mbox{\bf C}=\{C_{1},\ldots,C_{M}\}$, and select $1\leq K<M$ out of them. We proceed to define the two order-based rules for which we devise a secure MPC protocol in this section.
Copeland. Define for each $V_{n}$ a ballot matrix $P_{n}=(P_{n}(m,\allowbreak m^{\prime}):m,m^{\prime}\in[M])$, where $P_{n}(m,m^{\prime})=1$ if $C_{m}$ is ranked higher than $C_{m^{\prime}}$ in $V_{n}$’s ranking, $P_{n}(m,m^{\prime})=-1$ if $C_{m}$ is ranked lower than $C_{m^{\prime}}$, and all diagonal entries are 0. Then the aggregated ballot matrix,
@@
Maximin. Define the matrices $P_{n}$ so that $P_{n}(m,m^{\prime})=1$ if $C_{m}$ is ranked higher than $C_{m^{\prime}}$ in $V_{n}$’s ranking, while $P_{n}(m,m^{\prime})=0$ otherwise. As in Copeland rule, we let $P$ denote the sum of all ballot matrices, see Eq. (3). Then $P(m,m^{\prime})$ is the number of voters who preferred $C_{m}$ over $C_{m^{\prime}}$. The final score of $C_{m}$, $m\in[M]$, is then set to $\mbox{\bf w}(m):=\min_{m^{\prime}\neq m}P(m,m^{\prime})$.
3.2 Characterization of legal ballot matrices in the Copeland and Maximin rules
@@
    The set $\{Q_{m}:=\sum_{m^{\prime}\in[M]}Q(m^{\prime},m)\}_{m\in[M]}$ consists of $M$ distinct values.
@@
    The set $\{Q_{m}:=\sum_{m^{\prime}\in[M]}Q(m^{\prime},m)\}_{m\in[M]}$ consists of $M$ distinct values.
@@
$\Gamma(Q):=(Q(m,m^{\prime}):1\leq m<m^{\prime}\leq M)\,.$
Conditions 2 and 3 in Theorems 3 and 4 imply that every ballot matrix, $P_{n}$, is fully determined by its upper triangle, $\Gamma(P_{n})$, in either of the two voting rules that we consider.
3.3 A secure voting protocol
Here we present Protocol 1, a privacy-preserving implementation of the Copeland and Maximin order-based rules. The protocol computes, in a privacy-preserving manner, the winners in elections that are governed by those rules. It has two phases.
Phase 1 (Lines 1-8) is the voting phase. Here, each voter $V_{n}$, $n\in[N]:=\{1,\ldots,N\}$, constructs his ballot matrix, $P_{n}$ (Line 2). He then samples a random share-generating polynomial of degree $D^{\prime}-1$ for each of the $M(M-1)/2$ entries in $\Gamma(P_{n})$, where $\Gamma$ is the projection mapping defined in Section 3.2 (Lines 3-5). Finally, he sends to each tallier his relevant share in each of those entries, namely, the value of the corresponding share-generating polynomial at $x=d$, $d\in[D]$ (Lines 6-7). Following that, the talliers engage in an MPC sub-protocol to verify the legality of $V_{n}$’s ballot, without actually recovering that ballot (Line 8). We describe the validation sub-protocol in Section 3.4. Ballots that are found to be illegal are discarded. (In such cases it is possible to notify the voter that his ballot was rejected and allow him to resubmit it.)
Phase 2 of the protocol (Lines 9-11) is carried out after the voting phase had ended. First (Lines 9-10), each of the talliers, $T_{d}$, $d\in[D]$, computes his $D^{\prime}$-out-of-$D$ share, denoted $G_{d}(m,m^{\prime})$, in the $(m,m^{\prime})$th entry of the aggregated ballot matrix $P$, see Eq. (3), for all $1\leq m<m^{\prime}\leq M$. The heart of the protocol is in Line 11: here, the talliers engage in an MPC sub-protocol in order to find the indices of the $K$ winning candidates. We describe the details of those computations in Sections 3.6 and 3.7.
@@ 3.3 A secure voting protocol
2       Construct the ballot matrix, $P_{n}$, according to the selected indexing of candidates and the voting rule; forall $1\leq m<m^{\prime}\leq M$ do
3             Select uniformly at random $a_{n,m,m^{\prime},j}\in{\bf Z}_{p}$, $11\leq j\leq D^{\prime}-$ 1; Set $g_{n,m,m^{\prime}}(x)=P_{n}(m,m^{\prime})+\sum_{j=1}^{D^{\prime}-1}a_{n,m,m^{\prime},j}x^{j}$;
@@
5             Send to $T_{d}$ the set $\{n,m,m^{\prime},g_{n,m,m^{\prime}}(d):1\leq m<m^{\prime}\leq M\}$;
6      After all talliers receive their shares in $V_{n}$’s ballot, they engage in an MPC sub-protocol to check its legality;
7forall $T_{d}$, $d\in[D]$ do
8       Set $G_{d}(m,m^{\prime})=\sum_{n\in[N]}g_{n,m,m^{\prime}}(d)$, for all $1\leq m<m^{\prime}\leq M$;
$T_{1},\ldots,T_{D}$ find the indices of the $K$ winners and publish them; Output : The $K$ winning candidates from C.
@@ 3.4 Verifying the legality of the cast ballots
Voters may attempt to cheat by submitting illegal ballots in order to help their preferred candidate, or in order to sabotage the proper aggregation of the ballots. In real-life electronic elections, where voters typically cast their ballots on certified computers in voting centers, the chances of hacking such computers and tampering with the software are small. However, for full-proof security and as a countermeasure against dishonest voters that might manage to hack the voting system, we proceed to describe an MPC sub-protocol that enables the talliers to verify the legality of each ballot, even though those ballots remain hidden from them.
We note that in case a ballot is found to be illegal, the talliers may reconstruct it (by means of interpolation from the shares of $D^{\prime}$ talliers) and use the recovered ballot as proof of the voter’s dishonesty. The ability to construct such proofs, which could be used in legal actions against dishonest voters, might deter voters from attempting to cheat in the first place.
@@
We proceed to explain how the talliers can verify the legality of the cast ballots in each of the two order-based rules. That validation is based on the characterizations of legal ballots as provided in Theorems 3 and 4 for Copeland and Maximin, respectively. Note that the talliers need only to verify conditions 1 and 4; condition 2 needs no verification since the voters do not distribute shares in the diagonal entries, as those entries are known to be zero; and condition 3 needs no verification since the voters distribute shares only in the upper triangle and then the talliers use condition 3 in order to infer the lower triangle from the shared upper triangle.
@@
Verifying condition 1. Consider the shares that a voter distributed in $\Gamma(Q)$, where $Q$ is his ballot matrix. The talliers need to verify that each entry in the shared $\Gamma(Q)$ is either $1$ or $-1$ in Copeland, or either $1$ or 00 in Maximin. The verification is performed independently on each of the $M(M-1)/2$ entries of the shared upper triangle. A shared scalar $x$ is in $\{-1,1\}$ (resp. in $\{0,1\}$) iff $(x+1)\cdot(x-1)=0$ (resp. $x\cdot(x-1)=0$). Hence, the talliers use their shares in $x$ to compute the product $(x+1)\cdot(x-1)$ for Copeland or $x\cdot(x-1)$ for Maximin, as described in Section 2.3. If the computed product is zero for each of the $M(M-1)/2$ entries of $\Gamma(Q)$, then $\Gamma(Q)$ satisfies condition 1 in Theorem 3 (Copeland) or in Theorem 4 (Maximin). If, on the other hand, some of the multiplication gates issue a nonzero output, then the ballot will be rejected.
Verifying condition 4. First, we make the following observation. Let $x,y\in{\bf Z}_{p}$ be two values that are shared among the talliers. Denote by $x_{d}$ and $y_{d}$ the $D^{\prime}$-out-of-$D$ shares that $T_{d}$ has in $x$ and $y$, respectively. Then if $a,b\in{\bf Z}_{p}$ are two publicly known field elements, it is easy to see that $ax_{d}+by_{d}$ is a proper $D^{\prime}$-out-of-$D$ share in $ax+by$, $d\in[D]$. Also $a+x_{d}$ is a proper $D^{\prime}$-out-of-$D$ share in $a+x$.

### v7

@@ 3 A secure order-based voting protocol
In this section, we describe our method for securely computing the winners in two order-based voting rules, Copeland and Maximin. [^4] Our protocols are designed for scenarios in which it is required to issue just the identity of the winners. In scenarios where more information is needed, such as a ranking of all candidates, or even a score for each candidate, our protocols can easily output also that additional information. We begin with formal definitions in Section 3.1. Then, in Section 3.2, we characterize legal ballot matrices for each of the two rules. Such a characterization is an essential part of our method since the talliers need to verify, in an oblivious manner, that each cast ballot is indeed legal, and does not hide a malicious attempt to cheat or sabotage the elections. In other words, the talliers need to verify the legality of each cast ballot only through its secret shares, i.e., without getting access to the actual ballot. The characterization that we describe in Section 3.2 will be used later on to perform such an oblivious validation.
@@
Then, in Section 3.3 we introduce our secure voting protocol. The protocol description includes sub-protocols for validating the cast ballots and a sub-protocol for computing the final election results from all legal ballots. The validation sub-protocols are described in Sections 3.4 and 3.5. The sub-protocols for computing the election results are described in Sections 3.6 and 3.7 for Copeland and Maximin rules, respectively.
@@
We conclude this section with a discussion on how to set the size of the underlying field in which all computations take place (Section 3.8), and a discussion of the overall security of our protocol (Section 3.9).
@@ 3.1 Formal definitions
We consider a setting in which there are $N$ voters, $\mbox{\bf V}=\{V_{1},\ldots,V_{N}\}$, that wish to hold an election over $M$ candidates, $\mbox{\bf C}=\{C_{1},\ldots,C_{M}\}$. The output of the election is a ranking of C. We proceed to define the two order-based rules for which we devise a secure MPC protocol in this section.
$\bullet$ Copeland. Define for each $V_{n}$ a matrix $P_{n}=(P_{n}(m,\allowbreak m^{\prime}):m,m^{\prime}\in[M])$, where $P_{n}(m,m^{\prime})=1$ if $C_{m}$ is ranked higher than $C_{m^{\prime}}$ in $V_{n}$’s ranking, $P_{n}(m,m^{\prime})=-1$ if $C_{m}$ is ranked lower than $C_{m^{\prime}}$, and all diagonal entries are 0. Then the sum matrix,
@@
$\bullet$ Maximin. Define the matrices $P_{n}$ so that $P_{n}(m,m^{\prime})=1$ if $C_{m}$ is ranked higher than $C_{m^{\prime}}$ in $V_{n}$’s ranking, while $P_{n}(m,m^{\prime})=0$ otherwise. As in Copeland rule, we let $P$ denote the sum of all ballot matrices, see Eq. (1). Then $P(m,m^{\prime})$ is the number of voters who preferred $C_{m}$ over $C_{m^{\prime}}$. The final score of $C_{m}$, $m\in[M]$, is then set to $\mbox{\bf w}(m):=\min_{m^{\prime}\neq m}P(m,m^{\prime})$.
In both these order-based rules, we shall refer hereinafter to the matrix $P_{n}$ as $V_{n}$’s ballot matrix, and to $P$ as the aggregated ballot matrix.
3.2 Characterization of legal ballot matrices
@@
    If $Q_{m}:=\sum_{m^{\prime}\in[M]}Q(m,m^{\prime})$ then $Q_{m}\neq Q_{m^{\prime}}$ for all $1\leq m<m^{\prime}\leq M$.
@@
    If $Q_{m}:=\sum_{m^{\prime}\in[M]}Q(m,m^{\prime})$ then $Q_{m}\neq Q_{m^{\prime}}$ for all $1\leq m<m^{\prime}\leq M$.
Conditions 1-3 in both theorems are clear. As for condition 4, it states that the pairwise relations that $Q$ defines are consistent with a linear ordering of the $M$ candidates. As a counter-example, consider the following Copeland ballot matrix over $M=3$ candidates:
$Q=\left(\begin{array}[]{rrr}0&1&-1\\ -1&0&1\\ 1&-1&0\end{array}\right)\,.$
It complies with conditions 1-3 in Theorem 1, but violates condition 4 since $Q_{1}=Q_{2}=Q_{3}=0$. Hence, it is an illegal matrix. Indeed, the matrix states that $C_{1}>C_{2}$ (since $Q(1,2)=1$), $C_{2}>C_{3}$, and $C_{3}>C_{1}$. Such cyclicity is impossible in a linear ordering.
Proof of Theorem 1. Assume that the ordering of a voter over the set of candidates C is $(C_{j_{1}},\ldots,C_{j_{M}})$ where $\mbox{\bf j}:=(j_{1},\ldots,j_{M})$ is some permutation of $[M]$. Then the ballot matrix that such an ordering induces is
$Q=(Q(m,m^{\prime}))_{1\leq m,m^{\prime}\leq M}$
where $Q(m,m^{\prime})=1$ if $m$ appears before $m^{\prime}$ in the sequence j, $Q(m,m^{\prime})=-1$ if $m$ appears after $m^{\prime}$ in j, and $Q(m,m)=0$ for all $m\in[M]$. Such a matrix clearly satisfies conditions 1, 2 and 3 in the theorem. It also satisfies condition 4, as we proceed to show. Fix $m\in[M]$ and let $k\in[M]$ be the unique index for which $m=j_{k}$. Then the $m$th row in $Q$ consists of exactly $k-1$ entries that equal $-1$, $M-k$ entries that equal $1$, and a single entry on the diagonal that equals 0. Hence, $Q_{m}$, which is the sum of entries in that row, equals $M+1-2k$. Clearly, all those values are distinct, since the mapping $m\mapsto k$ is a bijection. That completes the first part of the proof: every legal Copeland ballot matrix satisfies conditions 1-4.
Assume next that $Q$ is an $M\times M$ matrix that satisfies conditions 1-4. Then for each $m\in[M]$, the $m$th row in the matrix consists of a single 0 entry on the diagonal where all other entries are either $1$ or $-1$. Assume that the number of $-1$ entries in the row equals $k(m)-1$, for some $k(m)\in[M]$, while the number of $1$ entries equals $M-k(m)$. Then the sum $Q_{m}$ of entries in that row equals $M+1-2k(m)$. As, by condition 4, all $Q_{m}$ values are distinct, then $k(m)\neq k(m^{\prime})$ when $m\neq m^{\prime}$. Stated otherwise, the sequence $\mbox{\bf k}:=(k(1),\ldots,k(M))$ is a permutation of $[M]$. Let $\mbox{\bf j}:=(j_{1},\ldots,j_{M})$ be the inverse permutation of k; i.e., for each $m\in[M]$, $j_{k(m)}=m$. Then it is easy to see that the matrix $Q$ is the Copeland ballot matrix that corresponds to the ordering j. That completes the second part of the proof: every matrix that satisfies conditions 1-4 is a legal Copeland ballot matrix.
The proof of Theorem 2 is similar to that of Theorem 1 and thus omitted.
@@
$\Gamma(Q):=(Q(m,m^{\prime}):1\leq m<m^{\prime}\leq M)\,.$ (3)
Conditions 2 and 3 in Theorems 1 and 2 imply that every ballot matrix, $P_{n}$, is fully determined by its upper triangle, $\Gamma(P_{n})$, in either of the two voting rules that we consider.
3.3 The protocol
Here we present Protocol 1, a privacy-preserving implementation of the Copeland and Maximin order-based rules. The protocol computes, in a privacy-preserving manner, the winners in elections that are governed by those rules. It has two phases. We begin with a bird’s-eye view of those two phases, and afterward, we provide a more detailed explanation.
Phase 1 (Lines 1-8) is the voting phase. In that phase, each voter $V_{n}$, $n\in[N]:=\{1,\ldots,N\}$, constructs his ballot matrix, $P_{n}$ (Line 2), and then creates and distributes to all talliers corresponding $D^{\prime}$-out-of-$D$ shares, with $D^{\prime}=\lfloor(D+1)/2\rfloor$, as described in Section 2.1 (Lines 3-7). Following that, the talliers jointly verify the legality of the shared ballot (Line 8). Phase 2 of the protocol (Lines 9-11) is carried out after the voting phase had ended. In that second phase, the talliers perform an MPC sub-protocol on the distributed shares in order to find the winning candidates, while remaining oblivious to the actual ballots.
After constructing his ballot matrix in Line 2, voter $V_{n}$, $n\in[N]$, samples a random share-generating polynomial of degree $D^{\prime}-1$ for each of the $M(M-1)/2$ entries in $\Gamma(P_{n})$, where $\Gamma$ is the projection mapping defined in Section 3.2 (Lines 3-5). Then, $V_{n}$ sends each tallier $T_{d}$ his relevant share of each of those entries, namely, the value of the corresponding share-generating polynomial at $x=d$, $d\in[D]$ (Lines 6-7).
In Line 8, the talliers engage in an MPC sub-protocol to verify the legality of $V_{n}$’s ballot, without actually recovering that ballot. There are two validation sub-protocols that need to be executed and we describe them in Sections 3.4 and 3.5. Ballots that are found to be illegal are discarded. (In such cases it is possible to notify the voter that his ballot was rejected and allow him to resubmit it.)
Phase 2 (Lines 9-11) takes place at the end of the voting period, after the voters had cast their ballots. First (Lines 9-10), each of the talliers, $T_{d}$, $d\in[D]$, computes his $D^{\prime}$-out-of-$D$ share, denoted $G_{d}(m,m^{\prime})$, in the $(m,m^{\prime})$th entry of the aggregated ballot matrix $P$, see Eq. (1), for all $1\leq m<m^{\prime}\leq M$. This computation follows from the linearity of the secret-sharing scheme. Indeed, as $g_{n,m,m^{\prime}}(d)$ is $T_{d}$’s share of $P_{n}(m,m^{\prime})$ in a $D^{\prime}$-out-of-$D$ Shamir’s secret sharing, then $G_{d}(m,m^{\prime})=\sum_{n\in[N]}g_{n,m,m^{\prime}}(d)$ is a $D^{\prime}$-out-of-$D$ Shamir’s secret share of $P(m,m^{\prime})=\sum_{n\in[N]}P_{n}(m,m^{\prime})$.
Then, in Line 11, the talliers engage in an MPC sub-protocol in order to find the indices of the $K$ winning candidates. How can the talliers do that when none of them actually holds the matrix $P$? We devote our discussion in Sections 3.6 and 3.7 to that interesting computational challenge.
@@ 3.3 The protocol
2       Construct the ballot matrix, $P_{n}$, according to the selected indexing of candidates and the voting rule
3      forall $1\leq m<m^{\prime}\leq M$ do
4             Select uniformly at random $a_{n,m,m^{\prime},j}\in{\bf Z}_{p}$, $11\leq j\leq D^{\prime}-$ 1
5            Set $g_{n,m,m^{\prime}}(x)=P_{n}(m,m^{\prime})+\sum_{j=1}^{D^{\prime}-1}a_{n,m,m^{\prime},j}x^{j}$
@@
7             Send to $T_{d}$ the set $\{(n,m,m^{\prime},g_{n,m,m^{\prime}}(d)):1\leq m<m^{\prime}\leq M\}$
8      After all talliers receive their shares of $V_{n}$’s ballot, they engage in an MPC sub-protocol to check its legality
9forall $T_{d}$, $d\in[D]$ do
10       Set $G_{d}(m,m^{\prime})=\sum_{n\in[N]}g_{n,m,m^{\prime}}(d)$, for all $1\leq m<m^{\prime}\leq M$
[…]
Output : The $K$ winning candidates from C.
@@ 3.4 Verifying the legality of the cast ballots
Voters may attempt to cheat by submitting illegal ballots in order to help their preferred candidate. For example, a dishonest voter may send the talliers shares of the matrix $cQ$, where $Q$ is his true ballot matrix and $c$ is an “inflating factor” greater than 1. If such a cheating attempt remains undetected, then that voter would manage to multiply his vote by a factor of $c$. The shares of the illegal “inflated” matrix are indistinguishable from the shares of the legal ballot matrix (unless, of course, a majority of $D^{\prime}$ talliers collude and recover the ballot -- a scenario that is impossible under our assumption that the talliers have an honest majority). Hence, it is necessary to devise a mechanism that would enable the talliers to check that the shares they received from each of the voters correspond to a legal ballot matrix, without actually recovering that matrix.
Malicious voters can sabotage the elections also in other manners. For example, a voter may create a legal ballot matrix $Q$ and then alter the sign of some of the ±$\pm 1$ entries in the shared upper triangle $\Gamma(Q)$, so that the resulting matrix no longer corresponds to an ordering of the candidates. Even though such a manipulated matrix cannot serve a dishonest voter in an attempt to help a specific candidate, it still must be detected and discarded. Failing to detect the illegality of such a ballot may result in an aggregated matrix $P$ that differs from the aggregated matrix $P^{\prime}$ that corresponds to the case in which that ballot is rejected. In such a case, the set of winners that $P$ would determine may differ from that which $P^{\prime}$ determines. In summary, it is essential to validate each of the cast ballots in order to prevent any undesirable sabotaging of the elections.
@@
In real-life electronic elections where voters typically cast their ballots on certified computers in voting centers, the chances of hacking such computers and tampering with the software that they run are small. However, for full-proof security and as a countermeasure against dishonest voters that might manage to hack the voting system, we proceed to describe an MPC solution that enables the talliers to verify the legality of each ballot, even though those ballots remain hidden from them. We note that in case a ballot is found to be illegal, the talliers may reconstruct it (by means of interpolation from the shares of $D^{\prime}$ talliers) and use the recovered ballot as proof of the voter’s dishonesty. The ability to construct such proofs, which could be used in legal proceedings against dishonest voters, might deter voters from attempting cheating in the first place.
@@
We proceed to explain how the talliers can verify the legality of the cast ballots in each of the two order-based rules. That validation is based on the characterizations of legal ballots as provided in Theorems 1 and 2 for Copeland and Maximin, respectively. Note that the talliers need only to verify conditions 1 and 4 (in either Theorem 1 or 2); condition 2 needs no verification since the voters do not distribute shares of the diagonal entries, as those entries are known to be zero; and condition 3 needs no verification since the voters distribute shares only in the upper triangle and then the talliers use condition 3 in order to infer the lower triangle from the shared upper triangle.
3.4.1 Verifying condition 1
@@
Consider the shares that a voter distributed in $\Gamma(Q)$, where $Q$ is his ballot matrix. The talliers need to verify that each entry in the shared $\Gamma(Q)$ is either $1$ or $-1$ in Copeland, or either $1$ or 00 in Maximin. The verification is performed independently on each of the $M(M-1)/2$ entries of the shared upper triangle. A shared scalar $x$ is in $\{-1,1\}$ (resp. in $\{0,1\}$) iff $(x+1)\cdot(x-1)=0$ (resp. $x\cdot(x-1)=0$). Hence, the talliers use their shares of $x$ in order to compute the product $(x+1)\cdot(x-1)$ for Copeland or $x\cdot(x-1)$ for Maximin, using the procedure for evaluating arithmetic functions (as described in Section 2.2 and further discussed in Section A.1 in the Appendix). If the computed results are zero for all $M(M-1)/2$ entries of $\Gamma(Q)$, then $\Gamma(Q)$ satisfies condition 1 in Theorem 1 (Copeland) or in Theorem 2 (Maximin). If, on the other hand, at least one of the results is nonzero, then the ballot will be rejected.
3.4.2 Verifying condition 4